OpenStudy (anonymous):

Doing a related rates problem to find the change in volume of a cone. I know volume of a cone is V=(1/3)bh. What would (dV/dt) equal?

OpenStudy (anonymous):

out of context, 0. the full problem would be required to answer properly

OpenStudy (anonymous):

To clarify, you know how you have to differentiate V before you can plug things in? I just want to know what that would be

OpenStudy (anonymous):

Ok. The radius r and height h of a circular cone change at a rate of 2 cm/s. How fast is the volume of the cone increasing when r=10 and h=20?

OpenStudy (anonymous):

\[\frac{ dV(t) }{ dt } = \frac{ 1 }{ 3 }\frac{ d }{ dt }(b(t)*h(t))\]

OpenStudy (anonymous):

and since we use product rule: \[\frac{ dV }{ dt } = \frac{ 1 }{ 3 }(\frac{ db }{ dt }h + b\frac{ dh}{dt })\]

OpenStudy (anonymous):

Thanks! I'm sorry not trying to be difficult but I suck at these problems, so why/how did you get that?

OpenStudy (anonymous):

Ohh product rule. ok gotcha!

OpenStudy (anonymous):

db/dt and dh/dt are given atually. hold on. this is b not r. so it's actually: \[\frac{ dV }{ dt } = \frac{ \pi }{ 3 }\frac{ d }{ dt }(r^2 * h)\]

OpenStudy (anonymous):

but yes. product rule lets you plug in r, h, dr/dt and dh/dt

OpenStudy (anonymous):

Where did pi come from? Different Equation?

OpenStudy (anonymous):

in V = 1/3 b*h, b is the base of the cone. and the base is the area of a circle: pi*r^2 V = (1/3)bh = (1/3)pi*r^2*h

OpenStudy (anonymous):

and we have to use r since we're given the rate of change of r and r's relation to b is not linear

OpenStudy (anonymous):

Oh right! Ok thanks that helped a lot!

OpenStudy (anonymous):

glad i could help. lmk if you want me to check your answer and if you have more questions :)

OpenStudy (anonymous):

I'm sorry, I got completely stuck so I ditched the problem and just came back to it. so given that last equation, would it come out to be: \[\frac{ dV }{ dt }= \frac{ \pi }{ 3 }2r ^{3}h\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Euler signed off, could anyone explain it to me from here reading our conversation?